OpenStudy (anonymous):

how do you solve this equation y=x-7 3x-3y=21

OpenStudy (p0sitr0n):

same equation? 3x-3y=21 x-y=7 -y=7-x y=x-7

OpenStudy (anonymous):

i dont understand what is the y

OpenStudy (p0sitr0n):

its a variable which depends on x actually, y=x-7 is a line